how would I solve this pre-alg. question ?
odina'a scores on the first three of four 100-pionts tests were 89, 92, and 82. Use the inequality 89+92+82+s/4 greater than or equal to 90, to determine what scores she could receive on the fourth test and have an average of at least 90 for all the tests

Answer by TimothyLamb(4379)

---
(89+92+82+s)/4 >= 90
89+92+82+s >= 90*4
263 + s >= 360
s >= 360 - 263
s >= 97
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php